Solve equation for all value of ø if ø measure in degrees
cos^2ø -5/2cosø -3/2=0

Answer by ewatrrr(10682)


Hi
cos^2ø -5/2cosø -3/2=0
factoring
(cosø -3)(cosø + 1/2)= 0
(cosø -3)= 0
cosø = 3
(cosø + 1/2)= 0
cosø = -1/2
ø is 120° or 240°